# A Decentralized Pilot Assignment Algorithm for Scalable O-RAN Cell-Free Massive MIMO

### v4

@@ III-E Iterative Codebook Search (CS) Algorithm
Suppose it is the turn of the $w$-th element of $\mathcal{U}^{(v)}_{\mathsf{cs}}$, denoted by $u_{v,w}$, to perform the isolated CS, where $w=1,2,\ldots,|\mathcal{U}^{(v)}_{\mathsf{cs}}|$. We first define $\ell_{v,w}$ to be the near-RT loop in which the agent $u_{v,w}$ begins its search. We also let $N_{\mathsf{s}}$ define the number of near-RT loops to be spent for codebook evaluation. During the first $N_{\mathsf{s}}$ near-RT loops (i.e., $\ell_{v,w}\leq\ell<\ell_{v,w}+N_{\mathsf{s}}$), the quality of current codebook matrix $\mathbf{T}^{\mathsf{old}}_{v,w}=[\bm{\phi}_{u_{v,w},1}^{(N_{\mathsf{n}}\ell_{v,w})},\bm{\phi}_{u_{v,w},2}^{(N_{\mathsf{n}}\ell_{v,w})},\ldots,\bm{\phi}_{u_{v,w},T_{p}}^{(N_{\mathsf{n}}\ell_{v,w})}]$ is evaluated by computing
@@
$\mathbf{T}_{u_{v,w}}^{(N_{\mathsf{n}}(\ell_{v,w}+2N_{\mathsf{s}}))}=\begin{cases}\mathbf{T}^{\mathsf{new}}_{v,w}&\text{if }\;\bar{r}^{\mathsf{new}}_{v,w}>\bar{r}^{\mathsf{old}}_{v,w},\\ \mathbf{T}^{\mathsf{old}}_{v,w}&\text{otherwise}.\end{cases}$ (29)
1 Input: Pilot length $T_{p}$, number of consistent PAs required for stability $N_{\mathsf{cs}}$, codebook evaluation interval $N_{\mathsf{s}}$, number of RT loops $N$, set of agents $\mathcal{U}$
2 Initialize CS run index $v=0$, set of agents subject for CS $\mathcal{U}^{(v)}_{\mathsf{cs}}=\varnothing$, the counter for agent $u$ $a_{u}=0$, $CS_{\mathsf{run}}=0$, and $CS_{\mathsf{iso}}=0$; assign identical codebook for all $u\in\mathcal{U}$; capture $\bm{S}^{(0)}_{u}$ using (13)
3 for $\ell=1$ to $N$ do
@@ 5             Capture $\bm{S}^{(\ell)}_{u}$ using (13)
As the CS described above runs for each agent in $\mathcal{U}^{(v)}_{\mathsf{cs}}$, total $2N_{\mathsf{s}}|\mathcal{U}^{(v)}_{\mathsf{cs}}|$ near-RT loops are spent to complete the CS run $v$. For every run, each agent tries a new codebook generated using a random rotation and decides to keep whichever codebook that yields higher reward. The algorithm starts its very first CS run at $\ell=\min_{u\in\mathcal{U}}\ell^{\star}_{u}$ and continuously conducts each subsequent CS run. By changing the codebook only when it is determined to be better, the algorithm proceeds to find the best set of $U$ codebooks that minimizes the degree of PC. Note that, in order to evaluate the codebooks, our CS scheme utilizes the reward $r^{(\ell)}_{u}(p)$, which is obtained during our PA scheme. Therefore, no additional information needs to be collected the O-DUs to conduct the CS. The overall procedure for our CS scheme is summarized in Alg 2.
@@ IV-C1 Comparison in MSE
Next, we consider scenarios in which users move over time (i.e., $\beta^{(i)}_{km}$ changes over $i$, and $v_{k}>0,\forall k\in\mathcal{K}$). Fig. 9 shows the sum-MSE performance (moving-averaged with a window size of $500$) of different PA algorithms with $K=24$ evaluated at three different user velocities: $1$, $5$, and $10$ km/h. The values of user velocity were selected so that the users still remain in the coverage area after their movement. PA solution obtained by the baselines at the beginning (i.e, $i=0$) becomes less effective as time advances, showing a different degree of steady increase by the velocity. Unlike the baselines, as our schemes make their decisions based on the real-time observations, in PA-DRL+MSG+CBS, PAs can be performed in an adaptive manner, maintaining its performance as shown in Fig. 9. Hence, our scheme can provide competitive performance with the prior knowledge-constrained baseline methods under a dynamic environment.
(a) Velocity = $1$ km/h
(b) Velocity = $5$ km/h
@@
Figure 9: MSE performance of different PA schemes over 24 mobile users with different velocities: $1$ km/h, $5$ km/h, and $10$ km/h.

### v5

@@ III-E Iterative Codebook Search (CS) Algorithm
[…]
@@
$\hskip-28.45274pt\mathbf{T}_{u_{v,w}}^{(N_{\mathsf{n}}(\ell_{v,w}+2N_{\mathsf{s}}))}=\begin{cases}\mathbf{T}^{\mathsf{new}}_{v,w}&\text{if }\;\bar{r}^{\mathsf{new}}_{v,w}>\bar{r}^{\mathsf{old}}_{v,w},\\ \mathbf{T}^{\mathsf{old}}_{v,w}&\text{otherwise}.\end{cases}$ (29)
1Input: Pilot length $T_{p}$, number of consistent PAs required for stability $N_{\mathsf{cs}}$, codebook evaluation interval $N_{\mathsf{s}}$, number of RT loops NN, set of agents $\mathcal{U}$
2 Initialize CS run index $v=0$, set of agents subject for CS $\mathcal{U}^{(v)}_{\mathsf{cs}}$$\varnothing$, the counter for agent $u$ $a_{u}=0$, $CS_{\mathsf{run}}=0$, and $CS_{\mathsf{iso}}=0$; assign identical codebook for all $u\in\mathcal{U}$; capture $\bm{S}^{(0)}_{u}$ using (13)
3 for $\ell=1$ to NN do
@@ 5     Capture $\bm{S}^{(\ell)}_{u}$ using (13)
[…]
@@ IV-C1 Comparison in MSE
Next, we consider scenarios in which users move over time (i.e., $\beta^{(i)}_{km}$ changes over $i$, and $v_{k}>0,\forall k\in\mathcal{K}$). Fig. 9 shows the sum-MSE performance (moving-averaged with a window size of $500$) of different PA algorithms with $K=24$ evaluated at three different user velocities: $1$, 55, and $10$ km/h. The values of user velocity were selected so that the users still remain in the coverage area after their movement. PA solution obtained by the baselines at the beginning (i.e, $i=0$) becomes less effective as time advances, showing a different degree of steady increase by the velocity. Unlike the baselines, as our schemes make their decisions based on the real-time observations, in PA-DRL+MSG+CBS, PAs can be performed in an adaptive manner, maintaining its performance as shown in Fig. 9. Hence, our scheme can provide competitive performance with the prior knowledge-constrained baseline methods under a dynamic environment.
(a) Velocity = 11 km/h
(b) Velocity = 55 km/h
@@
Figure 9: MSE performance of different PA schemes over 24 mobile users with different velocities: 11 km/h, 55 km/h, and $10$ km/h.